OpenStudy (anonymous):

Find all solutions to the equation. cos^2(x) + 2cos x + 1 = 0 Please help me!!!

OpenStudy (xapproachesinfinity):

what did you do so far?

OpenStudy (xapproachesinfinity):

think of how do you solve an equation like this \[x^2+2x+1=0\]

OpenStudy (xapproachesinfinity):

the principal is the same with trig

OpenStudy (anonymous):

would it be like.. (cos x + 1)^2 = 0 cos x = -1 therefore : x = π + 2kπ where k is any integer

OpenStudy (anonymous):

those question marks are supposed to be pi signs that I tried to get online but you see how it is ha

OpenStudy (xapproachesinfinity):

haha seems a good answer :)

OpenStudy (anonymous):

wait so it's really. pi+2kpi where k is any integer?

OpenStudy (xapproachesinfinity):

well yes :)

OpenStudy (xapproachesinfinity):

good job^_^

OpenStudy (anonymous):

Yayyy!! Thank you!!!

OpenStudy (xapproachesinfinity):

haha welcome ^_^